With a height of 60 in, Vincent was the shortest president of a particular club in the past century. The dlub presidents of the past century have a mean height of 61.3 in and a standard deviation of 1.2 in a. What is the positive difference between Vincents height and the mean? b. How many standard deviations is that the difference found in part (al? c. Convert Vincents height to a z score. d. Ifwe consider usual heights to be those that convert to z soores between -2 and 2, is Vincents height usual or unusual? The positive difference between Vincents height and the mean isc. b. The difference is? standard deviations. Type an integer or decimal rounded to two decimal places as needed.) c. The z score is Type an integer or decimal rounded to two decimal places as needed d. Is Vincenfs height usual or unusua? Usual O Unusual ar hed, Lear ets st ay Click to select your answer(s).
Answer #1

Solution

The dif ference between johnsons height and the mean = 60-61.3 =-1.3

Question wants Positive di f frence 1.3

------------------------------------------------------------------------

1.3 Standard deviation 1.08 1.2

------------------------------------------------------------------------

x-11 60-61.3 ~-score = 1.2

------------------------------------------------------------------------

Usual

The z - score is considered to be usual if the value is between-2 and 2